OpenStudy (megannicole51):

e=series of (1/n!) n=0 b=infinity ***My professor never taught us how to find upper and lower bounds of course and I believe there will be an exam question on it so help would be appreciated:D (a) Find a lower bound for e by evaluating the first five terms of the series (b) Show that (1/n!) < or equal to (1/2^(n-1)) for n> or equal to 1. (c) Find an upper bound for e using part (b)

OpenStudy (anonymous):

Well the terms in this series are strictly positive, which means that if you evaluate the first five terms, the full sum is going to be greater than what you just found, right?

OpenStudy (inkyvoyd):

For a, just evaluate the first five terms of the series algebraically, as @Jemurray3 said

OpenStudy (inkyvoyd):

well, less algebra and more arithmetic I guess

OpenStudy (megannicole51):

how would i do that? just plug in 1,2,3,4,5 for n?

OpenStudy (inkyvoyd):

actually, n=0,1,2,3,4

OpenStudy (megannicole51):

i know this might sound stupid but i forgot what the ! means....is it just if you have (1/5!) its going to be ((1)/(1*2*3*4*5))?

OpenStudy (megannicole51):

and oh yeah thank you!

OpenStudy (inkyvoyd):

yup

OpenStudy (inkyvoyd):

n!=n(n-1)(n-2)(n-2)...(1)

OpenStudy (anonymous):

i am thinking that induction is the way to go for the second part, but there may be an easier way to see it

OpenStudy (megannicole51):

"n!=n(n-1)(n-2)(n-2)...(1)" .....what?

OpenStudy (inkyvoyd):

uhh, just ignore that I guess

OpenStudy (megannicole51):

so its 1,1,(1/2),(1/6),(1/24)

OpenStudy (inkyvoyd):

for part b, factor the expressions in the denominator, kinda like @satellite73 would say you could use induction; and yes, 1,1,1/2,1/6,1/24

OpenStudy (megannicole51):

what do u mean factor the expressions?

OpenStudy (inkyvoyd):

"Show that (1/n!) < or equal to (1/2^(n-1)) for n> or equal to 1." take a general term in the series n, and factor the denominator for 1/n!, showing that 1/n! >1/2^(n-1)

OpenStudy (anonymous):

i guess induction is kind of overkill, but you do have an \(n\) there the second part asks you to show \(\frac{1}{n!}<\frac{1}{2^{n-1}}\) but this is more or less obvious, since if \(n>2\) you have \[\frac{1}{n!}=\frac{1}{n}\times\frac{ 1}{n-1}\times \frac{1}{n-2}\times ...\frac{1}{2}<\frac{1}{2}\times \frac{1}{2}\times ...\times \frac{1}{2}\]

OpenStudy (watchmath):

For \(n\geq 2\) \[ \underbrace{2\cdots2}_{n-1}\leq 1\cdot 2\cdots n \]

OpenStudy (anonymous):

so guess what your upper bound is?

OpenStudy (inkyvoyd):

I never liked geometric series

OpenStudy (megannicole51):

im not sure im still confused with b

OpenStudy (megannicole51):

i dont either lol

OpenStudy (anonymous):

lets to it for \(n=4\)

OpenStudy (megannicole51):

the (1/2)s are confusing me on what u wrote above...

OpenStudy (anonymous):

\(\frac{1}{4!}=\frac{1}{4\times 3\times 2\times 1}=\frac{1}{4}\times \frac{1}{3}\times \frac{1}{2}\times 1\)

OpenStudy (megannicole51):

ooooooh!

OpenStudy (anonymous):

in this case \(n=4\) and so \(n-1=3\)

OpenStudy (anonymous):

and it should be clear that \[\frac{1}{4}\times \frac{1}{3}\times \frac{1}{2}<\frac{1}{2}\times \frac{1}{2}\times \frac{1}{2}\] since all the denominators on the left are greater than or equal to the denominator of 2 on the right

OpenStudy (anonymous):

generalize to show that \(\frac{1}{n!}<\frac{1}{2^{n-1}}\)

OpenStudy (megannicole51):

okay so then how do i find the upper bound?

OpenStudy (anonymous):

not really much to generalize the denominators on the left get larger and larger, whilst the denominator on the right is always 2

OpenStudy (anonymous):

use part b since \(\frac{1}{n!}<\frac{1}{2^{n-1}}\) you know \[e=\sum_{n=1}^{\infty}\frac{1}{n!}<\sum_{n=0}^{\infty}\frac{1}{2^{n-1}}\] and the one on the right is your upper bound

OpenStudy (anonymous):

another typo damn!\[e=\sum_{n=0}^{\infty}\frac{1}{n!}<\sum_{n=1}^{\infty}\frac{1}{2^{n-1}}\]

OpenStudy (megannicole51):

how do u know thats my upper bound?

OpenStudy (megannicole51):

the one on the right is bigger?

OpenStudy (anonymous):

christ yet another typo you have to add one to the number on the right, since we are starting at \(1\)

OpenStudy (anonymous):

ok lets go back to part b

OpenStudy (anonymous):

you showed that except for the first 1 in the sum \[1+1+\frac{1}{2}+\frac{1}{3!}+...\]you have \(\frac{1}{n!}<\frac{1}{2^{n-1}}\)

OpenStudy (megannicole51):

is part b wrong?

OpenStudy (anonymous):

that means that the sum of the number on the left must be less that the sum of the numbers on the right, because each number on the left (except for the first 1, which we are not counting) is smaller than each number on the right

OpenStudy (anonymous):

no it is not wrong, i said "lets go back" because that is what we are going to use to do part c

OpenStudy (megannicole51):

oh okay:) proceed then...:)

OpenStudy (anonymous):

we are using this fact: if \(a_n\le b_n\) for all \(n\) then \[\sum a_n\le\sum b_n\] which should be more or less obvious right?

OpenStudy (megannicole51):

so if the sum on the left is less than the sum on the right then why is the sum on the left the upper bound?

OpenStudy (anonymous):

it is not, the sum on the right, the one you know how to compute, is the upper bound

OpenStudy (megannicole51):

okay then that makes WAY more sense!

OpenStudy (anonymous):

question says " Find an upper bound for e using part (b)"

OpenStudy (anonymous):

since you know how to compute \[\sum \frac{1}{2^n}\] this should now be easy, just don't forget to add 1 to your answer, because of the first 1 in \(1+1+\frac{1}{2}+\frac{1}{3!}+...\)

OpenStudy (anonymous):

you want to post the question here?

OpenStudy (megannicole51):

 what i did was An=(1/3^n+1) and Bn=(1/3^n) and did the lim of An/Bn and it equals 1